OpenStudy (anonymous):

I need to be able to solve these without a calculator if possible. f(x)=(log_2(x))/x^2 FIND: f(1/2), f(1), f(4). I have never done this before without a calculator so if you could show me the process, I would appreciate it. Thanks

OpenStudy (anonymous):

Okay, so for f(1/2), how can you rewrite 1/2 as a power of 2?

OpenStudy (anonymous):

I honestly have no idea

OpenStudy (anonymous):

So, you can rewrite 1/2 as 2^-1. Now, you have log_2(2^-1)/(1/2)^2. What's log_2(2^-1)?

OpenStudy (anonymous):

Okay that is pretty simple. I have never done this so most of it I probably won't be able to do.

OpenStudy (anonymous):

so have you got the first one yet?

OpenStudy (anonymous):

No. I don't know how to find the value of the log

OpenStudy (anonymous):

so here's a rule for you: \[\log_{a}a^b = b \]

OpenStudy (anonymous):

based on that, what is \[\log_{2} 2^{-1}\]

OpenStudy (anonymous):

-1

OpenStudy (anonymous):

right. So now we have \[f\left(\frac{1}{2}\right) =\frac{\log_{2}\left(2^{-1}\right)}{\left(\frac{1}{2}\right)^2}=-\frac{1}{\left(\frac{1}{4}\right)} = -4\]

OpenStudy (anonymous):

okay, i got it. The answer is negative. Does that work for all of them or is there different rules?

OpenStudy (anonymous):

no, lets try four. how do you rewrite 4 as a power of 2?

OpenStudy (anonymous):

well yeah the rule works for all of them but it's not always negative

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

2^2

OpenStudy (anonymous):

yep, so what's log_2(2^2)?

OpenStudy (anonymous):

1/8

OpenStudy (anonymous):

and f(1)=0

OpenStudy (anonymous):

yeah dude, you got it

OpenStudy (anonymous):

I have other log questions if you wouldnt mind helping me

OpenStudy (anonymous):

sure shoot

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

if log_a2=.36, log_a3=.56 and log_a=.83 fins an approximate value for each of the following a) log_a8 b) log_a(5/9) c)log_a12 d)log_a((fouthrt(2)/12)

OpenStudy (anonymous):

not fins find*

OpenStudy (anonymous):

what's the last one? I think you might have mistyped log_a=.83

OpenStudy (anonymous):

sorry log_a5=.83

OpenStudy (anonymous):

okay. so are you familiar with the rule \[\log a + \log b = \log ab\]?

OpenStudy (anonymous):

I have seen it but I have never used it

OpenStudy (anonymous):

okay. so based on that, what is log_a(2) + log_a(2) + log_a(2)?

OpenStudy (anonymous):

if you just multiply all the two's together

OpenStudy (anonymous):

*twos

OpenStudy (anonymous):

log_a6

OpenStudy (anonymous):

My mistake log_a8

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

so since we know log_a(2) = .36 and log_a(2)+log_a(2)+log_a(2) = log_a(8), what is the value of log_a(8)?

OpenStudy (anonymous):

1.08

OpenStudy (anonymous):

the second and fourth ones seem a little more complicated

OpenStudy (anonymous):

yeah it's the same deal for the second and fourth. another thing that might help is\[\log a -\log b = \log\left(\frac{a}{b}\right)\]

OpenStudy (anonymous):

Okay, I figured out the first three but I am lost with the fourth root

OpenStudy (anonymous):

would it be fourthroot(.36)-(.56+.83)=-1.3